OpenStudy (anonymous):

Carlos and Maria drove a total of 191 miles in 3.8 hours. Carlos drove the first part of the trip and averaged 49 miles per hour. Maria drove the remainder of the trip and averaged 52 miles per hour. For approximately hour many hours did maria drive? round your answer to the nearest tenth if necessary.

OpenStudy (anonymous):

help

OpenStudy (mathstudent55):

average speed = distance / time distance = average speed * time let t = time Maria drove then 3.8 - t = time Carlos drove distance Carlos drove = 49 * (3.8 - t) distance Maria drove = 52 * t total distance driven = 191 49(3.8 - t) + 52t = 191 Solve the equation for t.